# 00footnotetext: Shared first author; Shared last author

### v3

@@ Appendix H Hyperparameter study
A potential concern is that the training recipe can change conclusions about which dataset is optimal for training, due to interaction between training hyperparameters and dataset distributions. To address this confounder, we show in Table 13 that orderings between datasets are preserved for various combinations of weight decay and learning rate. Moreover we find that performance gains from optimal hyper-parameter choice and dataset design tend to be orthogonal and complement each other. We illustrate this effect in Table 14.
@@
Appendix I Model-based quality filters
We presented results for several different model-based quality filters in Section 4.4. In this section, we describe their implementations in further detail, focusing especially on fastText classifiers which were our method of choice for DCLM-baseline.
@@ Appendix I Model-based quality filters
We use the supervised fastText package from Joulin et al. [81] to train models to classify between chosen “high-quality” reference data which are given positive labels, and web-crawled data which are given negative labels. We then apply these classifiers to score each document from the pool we wish to filter, taking the predicted probability of the positive label as the score and computing a percentile-based threshold. In terms of training hyperparameters, we mostly used the default choices from the fastText package; the only hyperparameter change that we tried was to expand the feature space from unigrams only to both unigrams and bigrams (via setting the wordNgrams argument to be 2 instead of the default 1). This helped improve the quality of downstream filtered datasets, as shown in Table 15 (which extends Table 5 from Section 4.4).
@@
The bulk of our experimentation for training fastText models focused on constructing their underlying training sets, specifically the positively labeled reference data. For each experiment, we fixed the size of the training set to be 400K examples (i.e., 200K positive, 200K negative). The negatively labeled examples were sampled randomly from a set of documents that came from an earlier (smaller) version of our RefinedWeb reproduction. This version used trafilatura as the extractor instead of resiliparse, which we hypothesize might actually help for training the filtering model; as shown in Appendix J, trafilatura more aggressively removes boilerplate content that may appear in many pages (especiall from the same website). This type of content, if left in, may lead to the fastText models over-relying on these “spurious” features instead of the main contents of the page. For the positively labeled reference data, we tried several different sources, some of which involved further pre-processing:
@@
  • OH-2.5 + ELI5. Our goal for this mix was to source instruction and question-answer formatted data that is both high-quality and covers a wide range of potential topics. We sample 100K examples from OH-2.5, which we do not further pre-process. For ELI5, each raw page from the r/ExplainLikeImFive subreddit contains a post asking a specific question and then some number of comments aiming to answer said question. We curate examples for training fastText models by taking a post and combining it with the top-scoring answer (using the karma score derived from community up/down-votes). If there are ties, the longest answer is chosen. We also filter these examples by keeping only those where the post has score $\geq$ 3, the best comment has score $\geq$ 5, and there are at least 3 comments total.
@@
A recent line of work studies using instruction-tuned models as annotators to determine the potential usefulness of a document. Sachdeva et al. [139] proposed AskLLM, in which the authors prompted Flan-T5 models [37] to evaluate whether the given document $\dots$ contain[s] informative signal for pre-training a large-language model? An informative data point should be well-formatted, contain some usable knowledge of the world, and strictly NOT have any harmful, racist, sexist, etc. content.”. We implemented this method ourselves, testing several models as annotators, different settings for maximal sequence length, and several prompts on a small scale. We found that the best configuration was using Mistral-7B-Instruct-v0.2 [77], clipping the document at 1024 tokens, and taking the cumulative probabilities of Yes and yes tokens as the model score. We used the following prompt template:
         Evaluate the following paragraph for LLM pre-training suitability:      - Is it well-structured or contains useful examples to natural texts?      - Does it offer insights, useful facts or relevant information?      - Does it teach how to comply with open-ended tasks such as writing letters, poems, emails etc.?      - Is it free from harmful content?        If most criteria are met based on the content below, indicate ’yes’ for suitable. Otherwise, indicate ’no’ for unsuitable.        ### <input> ###        Is it suitable for LLM pre-training? OPTIONS:      - yes      - no
@@
Following the success of the different deduplication methods we used (Appendix K), we studied the effect of Semantic deduplication as proposed by Abbas et al. [1]. In this approach, the authors propose embedding the documents using pre-trained language models, clustering them using k-means, and removing all but one document from each group of closely related documents to encourage diversity in the dataset. We began by embedding each document in a pool of approximately 100 million documents (following Abbas et al. [1]’s best practices) with BGE-base [176]. We then used faiss-GPU [79] to perform spherical k-means clustering, with 20 iterations and $K=11000$. We sampled documents after discarding 25% of the data. As seen in Table 4, this intervention only negatively impacted the trained model. We hypothesize that the model used for embedding has a significant impact on the outcomes of this method. However, due to the large computational overhead when scaled, making it infeasible, we opted to rely on the deduplication methods outlined in Appendix K and leave this line of research for future work.
Appendix J Text extraction comparison
@@
We compute basic summary statistics for each extractor based on a sample of 10 WARC files (corresponding to 900K individual pages), presenting the results in Table 17. Notably, both resiliparse and trafilatura result at least 2x shorter documents on average compared to WET files. As shown in the examples in Section J.2, WET files indeed contain many additional lines with seemingly little value for pre-training (e.g. navigation bars, boilerplate notices, copyright statements). trafilatura and resiliparse trim most of these lines out, with the former being more strict about doing so. Between the two, resiliparse still keeps in about 10% more text; some of this additional text may provide useful content such as section titles and dates for articles. In terms of runtime, the two are much farther apart, with resiliparse being roughly 8x faster.
Table 17: Text extractor profiling. Characters and tokens are averaged over the number of resulting output pages (note that this may differ for each extractor due to due to the possibility of extraction failures). Throughput is measured in GBs of input WARCs processed per second for each CPU core.
Extractor Avg. Chars Avg. Tokens Throughput (GB / sec / core)
@@
MinHash is a locality-sensitive hashing technique used to group sets into collections based on their Jaccard similarity [28]. In the context of deduplicating text datasets, MinHash was first employed in Lee et al. [88] and then used in numerous other projects [121, 43]. We point readers to the main text of Lee et al. [88] and Appendix G.3.1 of Penedo et al. [121] for more details. The primary hyperparameters of note are the n-gram-size, and the number of permutations used. Following Lee et al. [88], Penedo et al. [121], we use an n-gram-size of 5 tokens and target a Jaccard similarity of 0.8. Departing from prior work, however, we modify the number of MinHash permutations used. Both Lee et al. [88] and Penedo et al. [121] use a total of 9,000 permutations, split into 450 buckets of 20 hashes each. We found this to be overly expensive and notice that similar Jaccard similarity plots can be attained with a much smaller number of permutations. For all of our ablations, we instead use a total of 1,395 permutations, split into 93 buckets of size 15. These hyperparameters were chose programmatically to mimic the Jaccard similarity plots as closely as possible, in an $\ell_{2}$ sense, with a fixed hash budget. See Figure 6 for more details.
@@
Bloom filters are a data structure that enable space-efficient set membership queries [26]. Explicitly, in sublinear space, a Bloom filter maintains a sketch of a set, that supports an insert operation, and a probabilistic membership_query operation, where the latter will never return any false negatives (i.e., return False for an element in the set), but will occasionally return a false positive (i.e., return True for an element not in the set). These were first used in the context of exact-duplicate removal in Soldaini et al. [150], but have since been extended to perform near-duplicate document and paragraph removal in a tool known as BFF (Big Friendly Filter) [67], and we further modify BFF to perform deduplication at the document and paragraph level simultaneously. We found that this technique is vastly more efficient than a MinhHash and SuffixArray pipeline. However there is one important caveat in that MinHash performs document-level deduplication at a document vs. document level, whereas BFF performs document-level deduplication at a document vs. corpus level.
@@
  • False Positive Rate: The two parameters that dictate the memory footprint required by BFF are the number of tokens and the false positive rate. However we only can control the false positive rate, and we notice that the Bloom filter size scales linearly with the negative log of the false positive rate. In particular, for a corpus of 1T tokens, occupying roughly 2TB of disk space, ensuring no false positives, i.e. setting the false positive rate to $1/1T$, would require 6.5TB of RAM. Here we argue analytically that a false positive rate of even as low as 0.01 suffices, which we support with experimentation in the next section.
@@
K.2.1 Deduplication ablations: pipeline at 1B-1x scale
We first perform ablations regarding the full pipeline choice for deduplication at the 1B-1x scale. We start with a pool of 76B tokens subsampled from Common Crawl with the preprocessing steps from Penedo et al. [121] applied. Then we apply a combination of deduplication steps, and subsample the pool further to the 28B tokens required for the 1B-1x scale. Finally we train and evaluate the Core score and the percentage of tokens that were removed by deduplication. The main questions we seek to answer from this round of ablations are:
@@
Table 18: Deduplication ablations (1B-1x scale). Starting from a pool of 76B tokens acquried from Common Crawl with the RefinedWeb Penedo et al. [121] pipeline applied, we evaluate the removal rate and Core score on different combinations of deduplication methods. Our Bloom filter method performs just as well as a combination of exact deduplication, MinHash and Suffix Array based techniques.
@@
76B 00% 40.1 +0.0
66B 13% 41.0 +0.9
62B 18% 40.9 +0.8
51B 33% 41.4 +1.3
56B 26% 41.7 +1.6
58B 24% 40.2 +0.1
49B 36% 41.3 +1.3
48B 37% 41.2 +1.2
45B 41% 41.7 +1.6
K.2.2 Deduplication ablations: pipeline at 7B-1x / 7B-2x scale
@@
For this round of ablations, we start with a pool sourced from one tenth of Common Crawl and run the preprocessing steps from Penedo et al. [121] and apply various deduplication pipelines. Then we subsample down to 138B tokens and train and evaluate models at the 7B-1x scale. The main questions we seek to answer from this round of ablations are:
@@
Results are contained in Table 19. The first point to note is that BFF with a min_ngram_size at 13 and 20 yields Core scores and MMLU scores that are comparable to the scores attained by a MinHash+SuffixArray deduplicated pool at the same scale. The second point to note regards the BFF min_ngram_size and sharding: interestingly a lower min_ngram_size yields higher Core scores, but lower MMLU scores. We also see that fewer shards decreases the token yield, but has variable effect on the Core score. We examine the hyperparameters for BFF more fully in the next subsection.
Table 19: Deduplication Ablations (7B-1x scale). Starting with a pool from Common Crawl and the RW-Filter pipeline processing applied, we compared several BFF hyperparameters against the MinHash and Suffix Array pipeline of [88, 121]. Our best BFF run and the prior works are bolded.
@@
Bloom Filter 5 13 32 25.0 52.0 4T
Bloom Filter 5 13 1 27.1 52.7 1.3T
Bloom Filter 13 13 32 28.7 51.7 3.8T
Bloom Filter 20 20 32 27.7 51.5 4T
Bloom Filter 20 20 10 28.4 51.2 3T
MinHash+SA N/A N/A 16 29.1 51.9 3.2T
Encouraged by these results, next we examine the top candidates for a scalable deduplication pipeline at the 7B-2x scale. Again we start with a pool obtained from one tenth of Common Crawl and generate several deduplicated pools. The questions of interest are the same as above and we summarize the results in Table 20. The key takeaways from this round of ablations is that at the 7B-2x scale, BFF with a min_ngram_size of 13 and 10 shards attains nearly identical performance to a MinHash+SuffixArray pipeline, whereas BFF with a min_ngram_size of 20 and 32 shards starts to lag behind, and that a min_ngram_size of 5 yields competitive Core scores, but falters in MMLU evaluations. While these experiments also vary the sharding choice, we view sharding primarily as a choice made to trade-off scalability with token yield. Larger shards are more expensive and less parallelizable and can decrease the token yield. For this round of ablations, the primary interest is to gain signal about how BFF compares to MinHash and Suffix Arrays at scale, and which are the correct hyperparameters for BFF. On this latter point, we chose to move forward with a min_ngram_size of 13 for generating DCLM-baseline.
Table 20: Deduplication Ablations (7B-2x scale). From the same pools as in Table 19, we trained and evaluated models at the 7B-2x scale. Notice that a min_ngram_size of 5 yields competitive Core results but drastically reduces MMLU scores.
@@
BFF 20 20 10 43.6 55.7 3T
BFF 13 13 10 44.3 55.5 3T
BFF 5 13 32 32.0 54.2 4T
MinHash+SA N/A N/A 16 44.4 55.4 3.2T
K.2.3 BFF hyperparameter ablations
While the above ablations largely focused on the Core score and MMLU as performance metrics, these are expensive and not suited for large swaths of ablations. Here we instead explore statistics of datasets deduplicated by BFF as we toggle the ngram_size hyperparameters, false positive rate, and input dataset size. We run separate experiments for each hyperparameter and finish each paragraph with the choice of hyperparameter we use for all larger scale runs.
@@
Here we start with the 75B token data pool as in Section K.2.1 and focus on a paragraph-only level BFF. In other words, we run BFF as described above, except omit the full-document removal step. We use the default hyperparameters for n-gram sizes as in Groeneveld [67], of 5 and 13 for min_ngram_size and max_ngram_size and a threshold of 0.8. We specifically look at the effect of changing the false positive rate and compute the removal rate (in bytes) of the output. From Table 21, we can see that a false positive rate of 0.1 suffices for a reasonably small pool such as this one. For larger pools, to be safe, we always set the false positive rate to 0.01.
@@
Table 22: BFF hyperparameter ablations. Starting with a pool of 341B tokens taken from Common Crawl with the RW-Filter pipeline applied, we run our Bloom filter deduplication with various hyperparameters noting how the document length and pool size change after deduplication. The input pool statistics are noted in the first row.
@@
K.2.4 Global MinHash on Open Datasets
@@
Appendix L Mixing sources
In Section 4.5, we showed that mixing our dataset with the usual sources did not improve its general performance, and hypothesized that this is due to the more stringent filtering performed for our Common Crawl portion. One could argue that improved filtering in the other sources could lead to similar improvements in performance. As such, we perform an experiment where we apply the same fastText classifier for filtering the other sources, as we do for our DCLM-baseline.
We take several source from RedPajama [160], and filter them with the fastText classifier applied in our DCLM-baseline, while keeping only the highest scored ones. We then add the resulting data to our pretraining dataset, and train models at the 1B-1x scale. The results of this can be seen in Table 25. We see that despite the more uniform handling of mixing across various sources, the additional sources still decrease performance.
We leave further analysis on potential mixtures of our datasets with other sources for future work. We also leverage the mixing track as a way for participants to explore such directions.
@@
Appendix M Human judgment
@@
We sampled 499 random documents from our pool, after going through the rule-based quality filters and deduplication. Figure 9 shows a histogram of the length of the documents in words. We asked 16 English speak AI graduate and professors to annotate each example as a good candidate to be included in an LM pretraining corpus (see instructions and some examples given to annotators below). [^6] All annotation efforts were done voluntarily and were not paid for. Out of the 499 samples, in 281 samples there was full agreement between all three annotators. We release both datasets in https://datacomp.ai/dclm.
@@
     Your task is to assess the quality of the documents (0-bad, 1-good), sampled randomly from RefinedWebV2, in terms of their usefulness for LLM training.    How do you judge the usefulness of a document?    While this is a subjective task, there are a few things to keep in mind:      1. We encourage you to review the details of the evaluation tasks that the LLMs are expected to excel in [link to spreadsheet with the tasks was provided]. Broadly, these tasks include reading comprehension, language understanding, world knowledge, and commonsense reasoning.      2. Check whether a particular document will be useful for answering any open-ended imperative tasks or questions. For example: writing a letter to a friend, writing a poem, replying to an email, cooking a dish, etc.      3. Check whether the data contains harmful content that you would not want an LLM to generate. This can include personal information, vulgar language, etc.      4. See the examples below to build some intuition.

### v4

@@ Appendix H Hyperparameter study
A potential concern is that differences in the training recipe can change conclusions about which dataset is optimal, due to interaction between training hyperparameters and dataset distributions. To address this confounder, we show in Table 12 that orderings between datasets are preserved for various combinations of weight decay and learning rate. Moreover we find that performance gains from optimal hyper-parameter choice and dataset design tend to be orthogonal and complement each other. We illustrate this effect in Table 13.
@@
Appendix I Architecture ablations
Figure 6: Performance of datasets in architecture variants versus our original architecture. We see that there is high correlation between the performance of a model trained with either a Gemma-like architecture or a Mamba-like architecture and the performance of the corresponding OpenLM one. This means that dataset improvements are consistent across model changes.
Similar to our hyperparameter study in Appendix H, here we explore whether our results also generalize across different architectures. We train models using two alternative architectures and examine the correlation of the results with the performance on the OpenLM architecture used in the rest of the paper. Specifically, we try the following:
  • First, we implement a Gemma [162] inspired variant of our base architecture, where we change the activation to GeGLU and include an RMS normalization layer. Here, we aim to test the effect of small changes to the architecture, while keeping the fundamental design the same (i.e., this variant is still a decoder-only, transformer based model).
  • Our second choice is the Mamba [74] architecture. Here, we aim to see how our datasets perform when the workings of the model change drastically (from a regular decoder-only model to a state-space model).
As shown in Figure 6, there is high correlation between the performance of a dataset when using the above two architecture variants and its performance for the original OpenLM architecture. This implies that the dataset improvements generalize across architectures.
Appendix J Model-based quality filters
We presented results for several different model-based quality filters in Section 4.4. In this section, we describe their implementations in further detail, focusing especially for fastText classifiers which were our method of choice for DCLM-baseline.
@@ Appendix I Architecture ablations
We use the fastText package from Joulin et al. [87] to train supervised models to classify between chosen “high-quality” reference data which are given positive labels, and web-crawled data which are given negative labels. We then apply these classifiers to score each document from the pool we wish to filter, taking the predicted probability of the positive label as the score and computing a percentile-based threshold. In terms of fastText’s training hyperparameters, we mostly used the default settings from the fastText package; the only hyperparameter change that we tried was to expand the feature space from unigrams only to both unigrams and bigrams (via setting the wordNgrams argument to be 2 instead of the default 1). This helped improve the quality of downstream filtered datasets, as shown in Table 14 (which extends Table 5 from Section 4.4).
@@
The bulk of our experimentation for training fastText models focused on constructing their underlying training sets, specifically the positively labeled reference data. For each experiment, we fixed the size of the training set to be 400K examples (i.e., 200K positive, 200K negative). The negatively labeled examples were sampled randomly from a set of documents that came from an earlier (smaller) version of our RefinedWeb reproduction. This version used trafilatura as the extractor instead of resiliparse, which we hypothesize might actually help for training the filtering model; as shown in Appendix K, trafilatura more aggressively removes boilerplate content that may appear in many pages (especiall from the same website). This type of content, if left in, may lead to the fastText models over-relying on these “spurious” features instead of the main contents of the page. For the positively labeled reference data, we tried several different sources, some of which involved further pre-processing:
@@
  • OH-2.5 + ELI5. Our goal for this mix was to source instruction and question-answer formatted data that is both high-quality and covers a wide range of potential topics. We sample 100K examples from OH-2.5, which we do not further pre-process. For ELI5, each raw page from the r/ExplainLikeImFive subreddit contains a post asking a specific question and then some number of comments aiming to answer said question. We curate examples for training fastText models by taking a post and combining it with the top-scoring answer (using the karma score derived from community up/down-votes). If there are ties, the longest answer is chosen. We also filter these examples by keeping only those where the post has score $\geq$ 0, the best comment has score $\geq$ 5, and there are at least 3 comments total.
@@
A recent line of work studies using instruction-tuned models as annotators to determine the potential usefulness of a document. Sachdeva et al. [146] proposed AskLLM, in which the authors prompted Flan-T5 models [40] to evaluate whether the given document $\dots$ contain[s] informative signal for pretraining a large-language model? An informative data point should be well-formatted, contain some usable knowledge of the world, and strictly NOT have any harmful, racist, sexist, etc. content.”. We implemented this method ourselves, testing several models as annotators, different settings for maximal sequence length, and several prompts on a small scale. We found that the best configuration was using Mistral-7B-Instruct-v0.2 [83], clipping the document at 1024 tokens, and taking the cumulative probabilities of Yes and yes tokens as the model score. We used the following prompt template:
         Evaluate the following paragraph for LLM pretraining suitability:      - Is it well-structured or contains useful examples to natural texts?      - Does it offer insights, useful facts or relevant information?      - Does it teach how to comply with open-ended tasks such as writing letters, poems, emails etc.?      - Is it free from harmful content?        If most criteria are met based on the content below, indicate ’yes’ for suitable. Otherwise, indicate ’no’ for unsuitable.        ### <input> ###        Is it suitable for LLM pretraining? OPTIONS:      - yes      - no
@@
Following the success of the different deduplication methods we used (Appendix L), we studied the effect of Semantic deduplication as proposed by Abbas et al. [1]. In this approach, the authors propose embedding the documents using pre-trained language models, clustering them using k-means, and removing all but one document from each group of closely related documents to encourage diversity in the dataset. We began by embedding each document in a pool of approximately 100 million documents (following Abbas et al. [1]’s best practices) with BGE-base [185]. We then used faiss-GPU [85] to perform spherical k-means clustering, with 20 iterations and $K=11000$. We sampled documents after discarding 25% of the data. As seen in Table 4, this intervention only negatively impacted the trained model. We hypothesize that the model used for embedding has a significant impact on the outcomes of this method. However, due to the large computational overhead when scaled, making it infeasible, we opted to rely on the deduplication methods outlined in Appendix L and leave this line of research for future work.
Appendix K Text extraction comparison
@@
We compute basic summary statistics for each extractor based on a sample of 10 WARC files (corresponding to 900K individual pages), presenting the results in Table 16. Notably, both resiliparse and trafilatura result in at least 2x shorter documents on average compared to WET files. As shown in the examples in Section K.2, WET files indeed contain many additional lines with seemingly little value for pretraining (e.g. navigation bars, boilerplate notices, copyright statements). trafilatura and resiliparse trim most of these lines out, with the former being more strict about doing so. Between the two, resiliparse still keeps in about 10% more text, some of which may provide useful content such as section titles and dates. In terms of runtime, the two are much farther apart, with resiliparse being roughly 8x faster.
Table 16: Text extractor profiling. Characters and tokens are averaged over the number of resulting output pages (note that this may differ for each extractor due to due to the possibility of extraction failures). Throughput is measured in MBs of input WARCs processed per second for each CPU core.
Extractor Avg. Chars Avg. Tokens Throughput (MB / sec / core)
resiliparse 3,227 1,329 4.55
trafilatura 2,901 1,179 0.56
@@
MinHash is a locality-sensitive hashing technique used to group sets into collections based on their Jaccard similarity [29]. In the context of deduplicating text datasets, MinHash has been employed to deduplicate GPT-3’s training dataset [31] and later an open-source implementation from the study of Lee et al. [94] was used in numerous other projects [128, 46]. We point readers to the main text of Lee et al. [94] and Appendix G.3.1 of Penedo et al. [128] for more details. The primary hyperparameters of note are the n-gram-size, and the number of permutations used. Following Lee et al. [94], Penedo et al. [128], we use an n-gram-size of 5 tokens and target a Jaccard similarity of 0.8. Departing from prior work, however, we modify the number of MinHash permutations used. Both Lee et al. [94] and Penedo et al. [128] use a total of 9,000 permutations, split into 450 buckets of 20 hashes each. We found this to be overly expensive and notice that similar Jaccard similarity plots can be attained with a much smaller number of permutations. For all of our ablations, we instead use a total of 1,395 permutations, split into 93 buckets of size 15. These hyperparameters were chosen programmatically to mimic the Jaccard similarity plots as closely as possible, in an $\ell_{2}$ sense, with a fixed hash budget. See Figure 7 for more details.
@@
Bloom filters are a data structure that enable space-efficient set membership queries [26]. Explicitly, in sublinear space, a Bloom filter maintains a sketch of a set, that supports an insert operation, and a probabilistic membership_query operation, where the latter will never return any false negatives (i.e., return False for an element in the set), but will occasionally return a false positive (i.e., return True for an element not in the set). These were first used in the context of exact-duplicate removal in Soldaini et al. [157], but have since been extended to perform near-duplicate document and paragraph removal in a tool known as BFF (Big Friendly Filter) [72], and we further modify BFF to perform deduplication at the document and paragraph level simultaneously. We found that this technique is vastly more efficient than a MinhHash and SuffixArray pipeline. However we make the important note that MinHash and BFF define duplicates in different ways: MinHash performs document-level deduplication at a document vs. document level, whereas BFF performs document-level deduplication at a document vs. corpus level.
@@
  • False Positive Rate: The two parameters that dictate the memory footprint required by BFF are the number of tokens and the false positive rate. However, we only can control the false positive rate, and we notice that the Bloom filter size scales linearly with the negative log of the false positive rate. In particular, for a corpus of 1T tokens, occupying roughly 2TB of disk space, ensuring no false positives, i.e. setting the false positive rate to $10^{-12}$, would require 6.5TB of RAM. Here we argue analytically that a false positive rate of even as low as 0.01 suffices, which we support with experimentation in the next section.
@@
L.2.1 Deduplication ablations: pipeline at 1B-1x scale
We first perform ablations regarding the full pipeline choice for deduplication at the 1B-1x scale. We start with a pool of 76B tokens subsampled from Common Crawl with the heuristic filtering steps from Penedo et al. [128] applied. Then we apply a combination of deduplication steps, and subsample the pool further to the 28B tokens required for the 1B-1x scale. Finally we train and evaluate the Core score and the percentage of tokens that were removed by deduplication. The main questions we seek to answer from this round of ablations are:
@@
Table 17: Deduplication ablations (1B-1x scale). Starting from a pool of 76B tokens acquried from Common Crawl with the RefinedWeb Penedo et al. [128] pipeline applied, we evaluate the removal rate and Core score on different combinations of deduplication methods. Our Bloom filter method performs as well as a combination of exact deduplication, MinHash and Suffix Array based techniques.
@@
76B 00% 24.7 +0.0
66B 13% 26.0 +1.3
62B 18% 25.6 +0.9
51B 33% 26.6 +1.9
56B 26% 26.8 +2.1
58B 24% 25.0 +0.3
49B 36% 26.2 +1.5
48B 37% 26.3 +1.6
45B 41% 26.8 +2.1
L.2.2 Deduplication ablations: pipeline at 7B-1x / 7B-2x scale
@@
For this round of ablations, we start with a pool sourced from one tenth of Common Crawl and run the preprocessing steps from Penedo et al. [128] and apply various deduplication pipelines. Then we filter using our fastText classifier and subsample down to 138B tokens to train and evaluate models at the 7B-1x scale. The main questions we seek to answer from this round of ablations are:
@@
Results are contained in Table 18. The first point to note is that BFF with min_ngram_size set to 13 and 20 yields Core scores and MMLU scores that are comparable to the scores attained by a MinHash+SuffixArray deduplicated pool at the same scale. The second point to note regards the BFF min_ngram_size and sharding: interestingly a lower min_ngram_size yields higher Core scores, but drastically lower MMLU scores if set to 5. We also see that fewer shards decreases the token yield, but has variable effect on the Core score. We examine the hyperparameters for BFF more fully in the next subsection.
Table 18: Deduplication Ablations (7B-1x scale). Starting with a tenth of DCLM-Pool filtered with RefinedWeb’s heuristic filters, we compare several BFF configurations against the MinHash and Suffix Array pipeline of [94, 128]. All datasets are then filtered with our fastText filter and token yields are projected based upon running the same pipeline on all of DCLM-Pool. Our best BFF run and the prior works are bolded.
@@
Bloom Filter 5 13 32 26.3 40.6 3.9T
Bloom Filter 5 13 1 27.1 41.7 1.3T
Bloom Filter 13 13 32 28.7 40.5 4.2T
Bloom Filter 20 20 32 27.7 40.0 4.3T
Bloom Filter 20 20 10 28.4 39.7 3.9T
MinHash+SA N/A N/A 16 29.1 40.8 3.2T
Encouraged by these results, next we examine the top candidates for a scalable deduplication pipeline at the 7B-2x scale. Again we start with a pool obtained from one tenth of Common Crawl and generate several deduplicated pools. The questions of interest are the same as above and we summarize the results in Table 19. The key takeaways from this round of ablations is that at the 7B-2x scale, BFF with a min_ngram_size of 13 and 10 shards attains nearly identical performance to a MinHash+SuffixArray pipeline, whereas BFF with a min_ngram_size of 20 and 32 shards starts to lag behind, and that a min_ngram_size of 5 yields competitive Core scores, but again falters on MMLU evaluations. While these experiments also vary the sharding choice, we view sharding primarily as a choice made to trade-off scalability with token yield. Larger shards are more expensive, less parallelizable, and can decrease the token yield. For this round of ablations, the primary interest is to gain signal about how BFF compares to MinHash and Suffix Arrays at scale, and which are the correct hyperparameters for BFF. On this latter point, we chose to move forward with a min_ngram_size of 13 for generating DCLM-baseline.
Table 19: Deduplication Ablations (7B-2x scale). Following the same setup as in Table 18, we now train and models at 7B-2x scale (280B tokens). Notice that a min_ngram_size of 5 again yields competitive Core results but drastically reduces MMLU scores.
@@
Bloom Filter 5 13 32 32.5 44.5 3.9T
Bloom Filter 13 13 10 44.3 45.3 3.8T
Bloom Filter 20 20 10 43.6 45.8 3.9T
MinHash+SA N/A N/A 16 44.4 45.5 3.2T
L.2.3 BFF hyperparameter ablations
While the above ablations largely focused on pre-training models and evaluating on Core and MMLU, running large swaths of ablations in this way remains computationally expensive. Here, we instead explore statistics of datasets deduplicated by BFF as we toggle the ngram_size hyperparameters, false positive rate, and input dataset size. We run separate experiments for each hyperparameter and finish each paragraph with the choice of hyperparameter we use for all larger scale runs.
@@
Here we start with the 75B token data pool as in Section L.2.1 and focus on a paragraph-only level BFF. In other words, we run BFF as described above, except omit the full-document removal step. We use the default hyperparameters for n-gram sizes as in Groeneveld [72], of 5 and 13 for min_ngram_size and max_ngram_size and a threshold of 0.8. We specifically look at the effect of changing the false positive rate and compute the removal rate (in bytes) of the output. From Table 20, we can see that a false positive rate of 0.1 suffices for a reasonably small pool such as this one. For larger pools, to be safe, we always set the false positive rate to 0.01.
@@
Table 21: BFF hyperparameter ablations. Starting with a pool of 341B tokens taken from DCLM-Pool and filtered with RefinedWeb’s heuristic filters, we run our Bloom filter deduplication with various hyperparameters noting how the document length and pool size change after deduplication. The input pool statistics are noted in the first row.
@@
L.2.4 Global MinHash on Open Datasets
@@
Appendix M Mixing sources
In Section 4.5, we showed that mixing our dataset with commonly used sources did not improve its general performance, and hypothesized that this is due to the more stringent filtering performed by DCLM-baseline on Common Crawl. Based on this hypothesis, one could argue that improved filtering of the other sources before mixing them with DCLM-baseline could also lead to improvements in performance. As one such attempt, we perform an experiment where we apply the same fastText classifier for filtering the other sources as we do for our DCLM-baseline.
We take several sources from RedPajama [168], and individually filter them with the fastText classifier applied in our DCLM-baseline, while keeping only the highest scored ones. We then add the resulting data to our pretraining dataset, and train models at the 1B-1x scale. The results of this can be seen in Table 24. We see that despite the more uniform handling of mixing across various sources, the additional sources still decrease performance.
We leave further analysis on potential filtering and mixing of non-Common Crawl sources for future work, hoping the mixing track can be used by participants to explore such directions.
@@
Appendix N Human judgment
@@
We sampled 499 random documents from our pool, after going through the rule-based quality filters and deduplication. Figure 10 shows a histogram of the length of the documents in words. We asked 16 English speaking AI graduate students and professors to annotate each example as a good candidate to be included in an LM pretraining corpus (see instructions and some examples given to annotators below). [^5] All annotation efforts were done voluntarily and were not paid for. Out of the 499 samples, in 281 samples there was full agreement between all three annotators. We release both datasets in our codebase.666 https://github.com/mlfoundations/dclm/tree/main/data
@@
     Your task is to assess the quality of the documents (0-bad, 1-good), sampled randomly from DCLM-RefinedWeb, in terms of their usefulness for LLM training.    How do you judge the usefulness of a document?    While this is a subjective task, there are a few things to keep in mind:      1. We encourage you to review the details of the evaluation tasks that the LLMs are expected to excel in [link to spreadsheet with the tasks was provided]. Broadly, these tasks include reading comprehension, language understanding, world knowledge, and commonsense reasoning.      2. Check whether a particular document will be useful for answering any open-ended imperative tasks or questions. For example: writing a letter to a friend, writing a poem, replying to an email, cooking a dish, etc.      3. Check whether the data contains harmful content that you would not want an LLM to generate. This can include personal information, vulgar language, etc.      4. See the examples below to build some intuition.